OpenStudy (firejay5):

Solve and show work! :D Medal Received if correct! :D 6. (1 over x - 1) | + (2 over x) = 0

OpenStudy (firejay5):

\[\frac{ 1 }{ x - 1 } + \frac{ 2 }{ x } = 0\]

OpenStudy (anonymous):

start with \[\frac{1}{x-1}=-\frac{2}{x}\] and then cross multiply